A pair of dice is thrown twice .Find the Probability of scoring 7 points

i) once ii) twice and iii) atleast once

Using binomial distribution


1
Expert's answer
2021-07-13T16:11:55-0400

The probability a roll of two dice will add to 7 is "\frac{6}{36}=\frac{1}{6}."

So, we have a binomial distribution with "n=2,\;p=\frac{1}{6}."

i) "P(X=1)=C_2^1(\frac{1}{6})^1(\frac{5}{6})^1=\frac{5}{18}."

ii) "P(X=2)=C_2^2(\frac{1}{6})^2=\frac{1}{36}."

iii) "P(X\ge1)=1-P(X=0)=1-C_2^0(\frac{5}{6})^2=\frac{11}{36}."